# Normality-Guided Distributional Reinforcement Learning for Continuous Control

### v3

@@ Abstract
Learning a predictive model of the mean return, or value function, plays a critical role in many reinforcement learning algorithms. Distributional reinforcement learning (DRL) has been shown to improve performance by modeling the value distribution, not just the mean. We study the value distribution in several continuous control tasks and find that the learned value distribution is empirical quite close to normal. We design a method that exploits this property, employ variances predicted from a variance network, along with returns, to analytically compute target quantile bars representing a normal for our distributional value function. In addition, we propose a policy update strategy based on the correctness as measured by structural characteristics of the value distribution not present in the standard value function. The approach we outline is compatible with many DRL structures. We use two representative on-policy algorithms, PPO and TRPO, as testbeds. Our method yields statistically significant improvements in 10 out of 16 continuous task settings, while utilizing a reduced number of weights and achieving faster training time compared to an ensemble-based method for quantifying value distribution uncertainty.
@@ 1 Introduction
In reinforcement learning, an agent receives rewards by interacting with the environment and updates their policy to maximize the cumulative reward or return. The return often has a high variance, which may make training unstable. To reduce the variance, actor-critic algorithms Thomas (2014); Schulman et al. (2015b); Mnih et al. (2016); Gu et al. (2016) use a learned value function, a model of the return that serves a baseline to reduce variance and speed up training.
@@
The standard approach to learning a value function estimates a single scalar value for each state. Several recent studies learn a value distribution instead, attempting to capture the randomness in the interaction between the agent and environment. One method for achieving this involves using multiple value functions (referred to as an ensemble) so that these functions collectively represent the value distribution Lee et al. (2020). An alternative method requires the value function to produce multiple outputs, represented as quantile bars Bellemare et al. (2017a); Barth-Maron et al. (2018); Dabney et al. (2017); Singh et al. (2020); Yue et al. (2020). These techniques, known as distributional reinforcement learning (DRL), have proven to enhance stability and expedite the learning process. The target for generating quantile bars is commonly derived from the distributional Bellman optimality operator, and the distributional value function is fitted to the target with the quantile regression Bellemare et al. (2017a); Dabney et al. (2017).
@@
The Markov chain central limit theorem (MC-CLT) suggests that in continuous tasks, the variance of the return keeps decreasing as the timestep goes. Moreover, when sufficient remaining timesteps exist, the return conforms to a normal distribution. We examine the value distribution trained using quantile regression and find, in continuous tasks, the predictions of the distribution closely resemble normal distributions, as described by the MC-CLT and depicted in Figure 1.
@@
In this paper, we assume that the value distribution of continuous tasks is governed by the MC-CLT rather than the distributional Bellman optimality operator, which prompts an increase in variance as the timestep progresses. To appropriately guide the distributional value function, we provide analytically computed quantile targets representing normal distributions, grounded in variances sourced from a variance network Nix and Weigend (1994); Kendall and Gal (2017) and returns. Our goal is to assess the accuracy of the current predictions made by the distributional value function. We aim to leverage distributional information, prioritizing values that are accurately estimated during policy updates.
@@
As we consider the reliable values for policy updates, we incorporate existing actor-critic algorithms Lee et al. (2020); Mai et al. (2022) that factor in uncertainty during policy updates as a comparison. These algorithms typically minimize the influence of samples with high variances. However, some states inherently possess elevated variances. For instance, the true variances are higher near the initial timestep because the expected return from each timestep gets smaller as timesteps goes on (Appendix D). The question arises: Should states having higher variances have less impact on policy updates? Although the predicted variance is correct for a state, its impact could be lessened simply because its variance is high. In contrast, we determine the impact of policy updates by evaluating how closely our distributional value function predicts the quantiles for normals. Given that our distributional value function is fitted to a normal quantile target, a well-trained distributional value function should yield quantiles that closely align with the shape of the normal distribution. This alignment indicates the function’s accuracy in reflecting its models’ underlying distribution. In Appendix E, we provide an additional discussion on tasks where the normality assumption is not applicable.
@@
  • We examine that the distributional value function approximates a normal for continuous tasks, and this normal’s variance increases with each timestep. To address this, we guide the distributional value function to produce quantiles representing a normal distribution, leveraging variances sourced from the variance network Nix and Weigend (1994); Kendall and Gal (2017) and sampled returns.
@@
  • We provide an empirical validation that uses PPO Schulman et al. (2017) and TRPO Schulman et al. (2015a) on several continuous control tasks. We compare our methods to standard algorithms (PPO and TRPO) as well as the ensemble-based approach. We find that our method exhibits better performance than the ensemble-based method in 14/16 tested environments, while using half as many weights and training twice as fast.
@@ 2 Related Work
We briefly review related work on distributional reinforcement learning (DRL) and supporting methods. Bellemare et al. Bellemare et al. (2017a) studied the distribution of the value network for deep Q-learning and proposed C51, a DRL algorithm that uses a categorical distribution to model the value distribution. C51 produced state-of-the-art performance on the Arcade Learning Environment Bellemare et al. (2013), but required the parameters of the modeling distribution to be fixed in advance. Dabney et al. Dabney et al. (2017) proposed QR-DQN, which leverages quantile regression and allows for an adaptive distribution structure. In addition, they resolved a key theoretical question with KL divergence ($D_{KL}$), which could lead to non-convergence, by replacing it with Wasserstein distance metric and showing convergence of the distributional value network. While these methods focus solely on the mean of the value distribution, our approach seeks to leverage the distributional characteristics through the normality assumption as well.
@@
Several extensions of DRL to continuous tasks have been proposed, including distributed distributional deterministic policy gradients (D4PG) Barth-Maron et al. (2018), sample-based distributional policy gradient (SDPG) Singh et al. (2020) and implicit distributional actor-critic (IDAC) Yue et al. (2020). Our method is model agnostic and compatible with any existing DRL algorithm that learns the value distribution with a single neural network.
@@
Compared to supervised learning, reinforcement learning has an unstable learning process, and uncertainty is often modeled to make RL learning more stable. We can decompose uncertainty into aleatoric and epistemic Hora (1996); Kiureghian and Ditlevsen (2009). Aleatoric comes from the stochasticity of data and is often measured with the attenuation loss function Nix and Weigend (1994). Epistemic uncertainty is caused by the limits of generalization of models and can be measured by dropout Kendall and Gal (2017) or ensemble methods Lakshminarayanan et al. (2016). The RL community models these two types of uncertainty to update the model safely or to support sparse reward functions Bellemare et al. (2016); Strehl and Littman (2005); Burda et al. (2018). Clements et al. Clements et al. (2019) penalizes actions with high aleatoric uncertainty, and epistemic uncertainty is used to guide exploration. To our knowledge, Moerland et al. Moerland et al. (2017) is the only previous work assuming a normal distribution for the return distribution, which they use to construct a bootstrap estimate of combined uncertainty. SUNRISE Lee et al. (2020) measures epistemic uncertainty with an ensemble technique and uses it to reduce the gradient of a sample with high uncertainty. Mai et al. Mai et al. (2022) leverage the two types of uncertainty with the combination of ensemble and attenuation loss. We utilize quantile bars, which represent the return’s distribution, as a proxy for epistemic uncertainty. This proxy is less resource-intensive in terms of training compared to ensemble methods, which significantly increase training costs (see Appendix of Mai et al. Mai et al. (2022)).
@@ 3.1 Reinforcement Learning
We consider an infinite-horizon Markov decision process (MDP) Puterman (2014); Sutton and Barto (2018) defined by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},{R},\gamma,\mu)$. The agent interacts with an environment and takes an action $a_{t}\in\mathcal{A}$ according to a policy $\pi_{\theta}(a_{t}|s_{t})$ parameterized by $\theta$ for each state $s_{t}\in\mathcal{S}$ at time $t$. Then, the environment changes the current state $s_{t}$ to the next state $s_{t+1}$ based on the transition probability $\mathcal{P}(s_{t+1}|s_{t},a_{t})$. The reward function ${R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ provides the reward for ($s_{t},a_{t}$). $\gamma$ denotes the discount factor and $\mu$ is the initial state distribution for $s_{0}$. The goal of a reinforcement learning (RL) algorithm is to find a policy $\pi_{\theta}$ that maximizes the expected cumulative reward, or return:
@@ 3.2 Distributional Reinforcement Learning
In distributional reinforcement learning (DRL), Bellemare et al. Bellemare et al. (2017b) introduced an action-state distributional value function $Q^{D,\pi}(s,a)$ which models the distribution of returns for an agent taking action $a$ and then following policy $\pi$. We model the distribution by having $Q^{D,\pi}$ output $N$ quantile bars $\{q_{0},q_{1},...,q_{N-1}\}$. Similar to the Bellman optimality operator $\mathcal{T^{\pi}}$ Watkins and Dayan (1992) (Equation 6) for the state-action value function $Q^{\pi}$, $Q^{D,\pi}$ is updated by a distributional Bellman optimality operator (Equation 7):
$\begin{split}&\mathcal{T^{\pi}}Q^{\pi}(s_{t},a_{t})=\mathbb{E}[R(s_{t},a_{t})]\\ &\qquad\quad\quad+\gamma\mathbb{E}_{s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})}[\max_{a_{t+1}}Q^{\pi}(s_{t+1},a_{t+1})]\\ &\text{where \,\,}s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t}),\\ &\qquad\quad a_{t+1}={\rm{argmax}_{a_{t+1}}}Q^{\pi}(s_{t+1},a_{t+1})\end{split}$ (6)
$\begin{split}&\mathcal{T^{\pi}}Q^{D,\pi}(s_{t},a_{t})=R(s_{t},a_{t})+\gamma Q^{D,\pi}(s_{t+1},a_{t+1})\\ &\text{where \,\,}s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t}),\\ &\qquad\quad a_{t+1}={\rm{argmax}_{a_{t+1}}}\mathbb{E}[Q^{D,\pi}(s_{t+1},a_{t+1})]\end{split}$ (7)
Figure 2: Illustration of how we measure uncertainty. When $N=4$, there are 4 quantile bars $\{q_{0},q_{1},q_{2},q_{3}\}$ as an output of $V^{D,\pi}$. Compute the mean $q_{\mathrm{{avg}}}$ of the quantile bars, and find $\sigma_{i}$ such that $P(Z\leq\frac{q_{i}-q_{\mathrm{{avg}}}}{\sigma_{i}})=0.2*(i+1)=\frac{i+1}{N+1}$ for $i\in\{0,1,2,3\}$. We consider $V^{D,\pi}$ as approximately $\mathcal{N}(q_{\mathrm{{avg}}},\sigma^{2}_{\mathrm{{avg}}})$ where $\sigma_{\mathrm{{avg}}}=\frac{1}{4}\sum_{i=0}^{3}\sigma_{i}$. We find the exact quantile location $\{q_{0}^{\prime},q_{1}^{\prime},q_{2}^{\prime},q_{3}^{\prime}\}$ of $\mathcal{N}(q_{\mathrm{{avg}}},\sigma^{2}_{\mathrm{{avg}}})$, then measure how much two types of quantiles are different with mean absolute error.
@@ 3.3 Markov Chain Central Limit Theorem
We introduce the Markov Chain Central Limit Theorem (MC-CLT) Jones (2004) (Theorem 1). We leverage this theorem to argue that returns are normally distributed in some continuous environments.
Theorem 1.
(Markov Chain Central Limit Theorem) Let $X$ be a Harris ergodic Markov chain on $X$ with stationary distribution $\rho$ and let $f:X\rightarrow\mathbb{R}$ is a Borel function. If $X$ is uniformly ergodic and $\mathbb{E}_{\rho}[f^{2}(x)]<\infty$ then for any initial distribution, as $n\rightarrow\infty$,
@@
To relate this to RL, we let $f$ be the reward function $R$ and $X$ be the state-action space. If the Markov chain induced by a policy is Harris ergodic, we expect that the theorem holds (as rewards being a Borel function is a mild condition). In detail, a return $G(s_{t})$ is consists of the sum of the rewards $\sum_{i=t}^{T}R(s_{t},a_{t})$. Thus, for sufficiently large $T$, the distribution of $G(s_{t})$ tends towards a normal distribution. This might not consistently be the case, particularly when $t$ closely approaches $T$ and $R(s_{t},\cdot)$ deviates from a normal for each possible action. Nevertheless, a distributional value function, trained using the Huber quantile regression, outputs a distribution that closely represents a normal for the final timestep, even for done states (Figure 1). It’s also worth noting that the presence of transient states can disrupt Harris ergodicity. However, this issue is not prevalent in many continuous control tasks, such as robotics environments in OpenAI Gym, since the optimal policy is stable and repetitive.

### v4

@@ Abstract
Learning a predictive model of the mean return, or value function, plays a critical role in many reinforcement learning algorithms. Distributional reinforcement learning (DRL) has been shown to improve performance by modeling the value distribution, not just the mean. We study the value distribution in several continuous control tasks and find that the learned value distribution is empirically quite close to normal. We design a method that exploits this property, employing variances predicted from a variance network, along with returns, to analytically compute target quantile bars representing a normal for our distributional value function. In addition, we propose a policy update strategy based on the correctness as measured by structural characteristics of the value distribution not present in the standard value function. The approach we outline is compatible with many DRL structures. We use two representative on-policy algorithms, PPO and TRPO, as testbeds. Our method yields statistically significant improvements in 10 out of 16 continuous task settings, while utilizing a reduced number of weights and achieving faster training time compared to an ensemble-based method for quantifying value distribution uncertainty.
@@ 1 Introduction
In reinforcement learning, an agent receives rewards by interacting with the environment and updates their policy to maximize the cumulative reward or return. The return often has a high variance, which may make training unstable. To reduce the variance, actor-critic algorithms (Thomas, 2014; Schulman et al., 2015b; Mnih et al., 2016; Gu et al., 2016) use a learned value function, a model of the return that serves a baseline to reduce variance and speed up training.
@@
The standard approach to learning a value function estimates a single scalar value for each state. Several recent studies learn a value distribution instead, attempting to capture the randomness in the interaction between the agent and environment. One method for achieving this involves using multiple value functions (referred to as an ensemble) so that these functions collectively represent the value distribution (Lee et al., 2020). An alternative method requires the value function to produce multiple outputs, represented as quantile bars (Bellemare et al., 2017a; Barth-Maron et al., 2018; Dabney et al., 2017; Singh et al., 2020; Yue et al., 2020). These techniques, known as distributional reinforcement learning (DRL), have proven to enhance stability and expedite the learning process. The target for generating quantile bars is commonly derived from the distributional Bellman optimality operator, and the distributional value function is fitted to the target with the quantile regression (Bellemare et al., 2017a; Dabney et al., 2017).
@@
Our goal is to derive quantile bars based on the Markov chain central limit theorem (MC-CLT) rather than the distributional Bellman optimality operator, and to introduce a novel uncertainty measure absent in standard distributional value functions, applying it to policy updates. The MC-CLT indicates that in continuous tasks, the variance of the return keeps decreasing as the timestep goes. Moreover, when sufficient remaining timesteps exist, the return conforms to a normal distribution. We examine the value distribution trained using quantile regression and find, in continuous tasks, the predictions of the distribution closely resemble normal distributions, as described by the MC-CLT and depicted in Figure 1. This contrasts with the distributional Bellman optimality operator, which causes the variance of the value distribution to increase over timesteps (Figure 2), due to the multiplication of $\gamma<1$ with the next state and action Q value (Equation LABEL:distributional_bellman_operator). Section 4.2 provides a detailed example of this variance increase.
@@
In this paper, we assume that the value distribution of continuous tasks is governed by the MC-CLT rather than the distributional Bellman optimality operator. To appropriately guide the distributional value function, we provide analytically computed quantile targets representing normal distributions, grounded in variances sourced from a variance network (Nix & Weigend, 1994; Kendall & Gal, 2017) and returns. We assess the accuracy of the current predictions made by the distributional value function. We aim to leverage distributional information, prioritizing values that are accurately estimated during policy updates.
@@
As we consider the reliable values for policy updates, we incorporate existing actor-critic algorithms (Lee et al., 2020; Mai et al., 2022) that factor in uncertainty during policy updates as a comparison. These algorithms typically minimize the influence of samples with high variances. However, some states inherently possess elevated variances. For instance, the true variances are higher near the initial timestep because the expected return from each timestep gets smaller as timesteps goes on (Appendix D). The question arises: Should states having higher variances have less impact on policy updates? Although the predicted variance is correct for a state, its impact could be lessened simply because its variance is high. In contrast, we determine the impact of policy updates by evaluating how closely our distributional value function predicts the quantiles for normals. Given that our distributional value function is fitted to a normal quantile target, a well-trained distributional value function should yield quantiles that closely align with the shape of the normal distribution. This alignment indicates the function’s accuracy in reflecting its models’ underlying distribution. In Appendix E, we provide additional discussion on tasks where the normality assumption may not be applicable, as well as experiments in the opposite direction, where samples that are not closely aligned are given higher weights for exploration.
@@
  • We examine that the distributional value function approximates a normal for continuous tasks, and this normal’s variance increases with each timestep. To address this, we guide the distributional value function to produce quantiles representing a normal distribution, leveraging variances sourced from the variance network (Nix & Weigend, 1994; Kendall & Gal, 2017) and sampled returns.
@@
  • We provide an empirical validation that uses PPO (Schulman et al., 2017) and TRPO (Schulman et al., 2015a) on several continuous control tasks. We compare our methods to standard algorithms (PPO and TRPO) as well as the ensemble-based approach. We find that our method exhibits better performance than the ensemble-based method in 10/16 tested environments, while using half as many weights and training twice as fast.
@@ 2 Related Work
We briefly review related work on distributional reinforcement learning (DRL) and supporting methods. Bellemare et al. (2017a) studied the distribution of the value network for deep Q-learning and proposed C51, a DRL algorithm that uses a categorical distribution to model the value distribution. C51 produced state-of-the-art performance on the Arcade Learning Environment (Bellemare et al., 2013), but required the parameters of the modeling distribution to be fixed in advance. Dabney et al. (2017) proposed QR-DQN, which leverages quantile regression and allows for an adaptive distribution structure. In addition, they resolved a key theoretical question with KL divergence ($D_{KL}$), which could lead to non-convergence, by replacing it with Wasserstein distance metric and showing convergence of the distributional value network. While these methods focus solely on the mean of the value distribution, our approach seeks to leverage the distributional characteristics through the normality assumption as well.
@@
Several extensions of DRL to continuous tasks have been proposed, including distributed distributional deterministic policy gradients (D4PG) (Barth-Maron et al., 2018), sample-based distributional policy gradient (SDPG) (Singh et al., 2020) and implicit distributional actor-critic (IDAC) (Yue et al., 2020). Our method is model agnostic and compatible with any existing DRL algorithm that learns the value distribution with a single neural network.
@@
Compared to supervised learning, reinforcement learning has an unstable learning process, and uncertainty is often modeled to make RL learning more stable. We can decompose uncertainty into aleatoric and epistemic (Hora, 1996; Kiureghian & Ditlevsen, 2009). Aleatoric comes from the stochasticity of data and is often measured with the attenuation loss function (Nix & Weigend, 1994). Epistemic uncertainty is caused by the limits of generalization of models and can be measured by dropout (Kendall & Gal, 2017) or ensemble methods (Lakshminarayanan et al., 2016). The RL community models these two types of uncertainty to update the model safely or to support sparse reward functions (Bellemare et al., 2016; Strehl & Littman, 2005; Burda et al., 2018). Clements et al. (2019) penalizes actions with high aleatoric uncertainty, and epistemic uncertainty is used to guide exploration. To our knowledge, Moerland et al. (2017) is the only previous work assuming a normal distribution for the return distribution, which they use to construct a bootstrap estimate of combined uncertainty. SUNRISE (Lee et al., 2020) measures epistemic uncertainty with an ensemble technique and uses it to reduce the gradient of a sample with high uncertainty. Mai et al. (2022) leverage the two types of uncertainty with the combination of ensemble and attenuation loss. We utilize quantile bars, which represent the return’s distribution, as a proxy for epistemic uncertainty. This proxy is less resource-intensive in terms of training compared to ensemble methods, which significantly increase training costs (see Appendix of Mai et al. (2022)).
@@ 3.1 Reinforcement Learning
We consider an infinite-horizon Markov decision process (MDP) (Puterman, 2014; Sutton & Barto, 2018) defined by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},{R},\gamma,\mu)$. The agent interacts with an environment and takes an action $a_{t}\in\mathcal{A}$ according to a policy $\pi_{\theta}(a_{t}|s_{t})$ parameterized by $\theta$ for each state $s_{t}\in\mathcal{S}$ at time $t$. Then, the environment changes the current state $s_{t}$ to the next state $s_{t+1}$ based on the transition probability $\mathcal{P}(s_{t+1}|s_{t},a_{t})$. The reward function ${R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ provides the reward for ($s_{t},a_{t}$). $\gamma$ denotes the discount factor and $\mu$ is the initial state distribution for $s_{0}$. The goal of a reinforcement learning (RL) algorithm is to find a policy $\pi_{\theta}$ that maximizes the expected cumulative reward, or return:
@@ 3.2 Distributional Reinforcement Learning
In distributional reinforcement learning (DRL), Bellemare et al. (2017b) introduced an action-state distributional value function $Q^{D,\pi}(s,a)$ which models the distribution of returns for an agent taking action $a$ and then following policy $\pi$. We model the distribution by having $Q^{D,\pi}$ output $N$ quantile bars $\{q_{0},q_{1},...,q_{N-1}\}$. Similar to the Bellman optimality operator $\mathcal{T^{\pi}}$ (Watkins & Dayan, 1992) (Equation LABEL:bellman_operator) for the state-action value function $Q^{\pi}$, $Q^{D,\pi}$ is updated by a distributional Bellman optimality operator (Equation LABEL:distributional_bellman_operator):
$\begin{split}&\mathcal{T^{\pi}}Q^{\pi}(s_{t},a_{t})=\mathbb{E}[R(s_{t},a_{t})]+\gamma\mathbb{E}_{s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})}[\max_{a_{t+1}}Q^{\pi}(s_{t+1},a_{t+1})]\\ &\text{where \,\,}s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t}),a_{t+1}={\rm{argmax}_{a_{t+1}}}Q^{\pi}(s_{t+1},a_{t+1})\end{split}$ (3)
$\begin{split}&\mathcal{T^{\pi}}Q^{D,\pi}(s_{t},a_{t})=R(s_{t},a_{t})+\gamma Q^{D,\pi}(s_{t+1},a_{t+1})\\ &\text{where \,\,}s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t}),a_{t+1}={\rm{argmax}_{a_{t+1}}}\mathbb{E}[Q^{D,\pi}(s_{t+1},a_{t+1})]\end{split}$ (4)
@@ 3.3 Markov Chain Central Limit Theorem
We introduce the Markov Chain Central Limit Theorem (MC-CLT) (Jones, 2004) (Theorem 1). We leverage this theorem to argue that returns are normally distributed in some continuous environments.
Theorem 1 (Markov Chain Central Limit Theorem).
Let $X$ be a Harris ergodic Markov chain on $X$ with stationary distribution $\rho$ and let $f:X\rightarrow\mathbb{R}$ is a Borel function. If $X$ is uniformly ergodic and $\mathbb{E}_{\rho}[f^{2}(x)]<\infty$ then for any initial distribution, as $n\rightarrow\infty$,
@@
To relate this to RL, we let $f$ be the reward function $R$ and $X$ be the state-action space. If the Markov chain induced by a policy is Harris ergodic, we expect that the theorem holds (as rewards being a Borel function is a mild condition). In detail, a return $G(s_{t})$ is consists of the sum of the rewards $\sum_{i=t}^{T}R(s_{t},a_{t})$. Thus, for sufficiently large $T$, the distribution of $G(s_{t})$ tends towards a normal distribution. This might not consistently be the case, particularly when $t$ closely approaches $T$ and $R(s_{t},\cdot)$ deviates from a normal for each possible action. Nevertheless, a distributional value function, trained using the Huber quantile regression, outputs a distribution that closely represents a normal for the final timestep, even for terminal states (Figure 1). It’s also worth noting that the presence of transient states can disrupt Harris ergodicity. However, this issue is not prevalent in many continuous control tasks, such as robotics environments in OpenAI Gym, since the optimal policy is stable and repetitive.
Figure 2: The standard deviation computed from each of the distributional value functions trained on the distributional Bellman optimality operator (Equation LABEL:distributional_bellman_operator) with exponential smoothing. The estimated standard deviation increases as timestep increases.